MILLIMETER-WAVE SPECTRUM OF THE NO DIMER

MATTHEW D. BROOKES and \underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; T. AMANO, Department of Physics and Astronomy, Ibaraki University, Mito 310, Japan.

We are studying the pure rotational spectrum of the NO dimer, (NO)_{2}. The dimers are produced using a continuous supersonic jet source, and probed using a mm-wave spectrometer based on harmonic multiplication. Microwave radiation (75-100 GHz) from a Gunn oscillator is multiplied (×3 or ×4) in a commercial tripler, focussed through the jet by teflon lenses, and detected by a helium-cooled InSb bolometer. The frequency scan and data collection are under microcomputer control. The computer also turns the jet off and on at intervals of 5 scans (about 5 sec), and automatically subtracts the background and sample spectra. Previously, 4 rotational transitions of (NO)_{2} in the 0 to 23 GHz microwave region had been reported.^{a} Currently, we have measured about 80 new transitions in the frequency range from 227 to 382 GHz, with J values from 5 to 16, and K_{a} values from 2 to 8. The underlying ^{14}N hyperfine structure is partially resolved for some of the transitions. The observed line positions are quite well predicted by the existing molecular parameters from our analysis of the ν_{1} infrared band,^{b} but the new data will obviously allow a considerably more precise set of parameters to be derived. These new mm-wave results do not directly address the two great mysteries of the NO dimer, namely the locations of the intermolecular vibrational modes and of the low-lying electronic states. We still hope to address the former mystery by means of long-path, low temperature FT spectroscopy in the far-ir region.

aC.M. Western, P.R.R. Langridge-Smith, B.J. Howard, and S.E. Novick, Mol. Phys. 44, 145 (1981); S. Kukolich, J. Mol.

Spectrosc. 98, 80 (1983).

bB.J. Howard and A.R.W. McKellar, Mol. Phys. 78, 55 (1993).